# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

### v1

@@ 3.6 Performance evaluation
The main advantage of GMASK is providing a fast response to queries at the expense of returning approximate results. Thus, a straightforward strategy to assess the performance of this algorithm is to undertake an exhaustive search of all elements in a given dataset $X$. Then, the proportion of data points that were correctly retrieved can be determined. This is the common approach to all evaluation experiments developed in this study.
A comprehensive series of experiments are conducted across datasets of diverse types, sizes, dimensionality, and sparsity to evaluate the performance of GMASK, which introduces $k$-medoids for the bottom-top construction of the index tree. This novelty enables its parametrisation with several $\ell_{p}$ norms as outlined in “k-medoids as a generalisation of k-means”. These results will be compared with those obtained using other state-of-the-art algorithms. As mentioned above, our objective is to understand better the impact of utilising distances other than the Euclidean metric. Through these experiments, we aim to demonstrate that data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
@@ 3.7 Benchmark algorithms for similarity search
To evaluate our thesis, we compare GMASK performance in approximate nearest neighbour search with respect to alternative MAM.
3.7.1 KD-Tree
The KD-Tree (K-Dimensional Tree) [40] is an exact similarity search method used to find the nearest neighbours to a given query point. It constructs a binary tree structure by recursively dividing the data based on the maximum spread along different dimensions and the search algorithm efficiently narrows down the search space by traversing the tree. In low-dimensional spaces, KD-trees can achieve sublinear search times. However, as the dimensionality increases, the performance of the KD-tree can degrade due to the Curse of Dimensionality phenomenon.
As it performs an exact search, the solutions provided by this method are used as a baseline to evaluate the performance of other approximate algorithms. We shall compare the k-nearest neighbours obtained by each method against those obtained using the KD-Tree to assess their accuracy.
3.7.2 FLANN
FLANN (Fast Library for Approximate Nearest Neighbours) [11] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets, extensively used in machine learning, computer vision, and data mining. FLANN employs techniques such as random projection and hierarchical subdivision to construct an index structure that enables faster search operations. It offers several index types, including KD-tree, $k$-means tree, and hierarchical clustering tree, which are suitable for diverse data types and query requirements. While FLANN has implementations in several programming languages, we use the Python bindings provided by the pyflann  [^2] https://github.com/primetang/pyflann library.
3.7.3 PyNNDescent
The PyNNDescent algorithm is an approximate similarity search method [41] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/. It is based on the NNDescent algorithm and used for efficient nearest neighbour search in large-scale datasets. PyNNDescent constructs a graph-based index by connecting each data point to its approximate nearest neighbours. It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching, and pruning. The algorithm also provides parameters to control the trade-off between search accuracy and computational cost. PyNNDescent is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining, and information retrieval tasks.
@@ 3.8 Datasets
Several experiments have been conducted to evaluate the performance of GMASK, using four datasets diverse in nature and public access. Table 1 briefly summarises their main features.
3.8.1 Municipalities
@@
Although this is a two-dimensional dataset with a low dispersion, it is an exceptionally intriguing case due to its intuitive representation and the presence of outliers, such as the Canary Islands and the Balearic Islands. The location of these islands deviates from the rest of the points significantly and poses a challenge to any similarity search method applied to this dataset.
3.8.2 MNIST
The MNIST_784 dataset  [^5] http://yann.lecun.com/exdb/mnist/ is a widely used machine learning benchmark dataset whose simplicity makes it an ideal starting point for algorithm development and evaluation. Firstly, it consists of 60,000 training images and 10,000 test images, a relatively small size that makes possible to carry out computationally efficient experimentation and prototyping. Secondly, it comprises grayscale images of handwritten digits ranging from 0 to 9, making it visually straightforward and easy to work with. Moreover, each image is 28x28 pixels in size, resulting in a total of 784 dimensions when flattened, allowing researchers to explore algorithms in the context of high-dimensional data. Many well-established techniques and algorithms have been evaluated on MNIST [42, 43], making it a reference point for evaluating the performance of approximate similarity search methods [16, 44].
3.8.3 GLOVE
@@
3.8.4 NYtimes
The NYtimes dataset, part of the Bags of Words dataset from the UCI repository [^7] https://archive.ics.uci.edu/dataset/164/bag+of+words, comprises a collection of New York Times news articles represented as a bag of words. Each document in the dataset is associated with a set of word occurrences, where the dimensions represent unique words extracted from the articles. The dataset is organised as a document–word matrix, where each row corresponds to a document and each column corresponds to a word. The values in the matrix indicate the frequency of each word occurring in the respective document. Preprocessing steps include tokenization, removal of stopwords, and vocabulary truncation, with only words occurring more than ten times retained.
The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version, and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [45].
Dataset Label n Dimensionality High Sparsity Data Type
Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial
MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image
GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text
NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text
@@ 4 Experimental results
In order to show the versatility of GMASK, we present the analysis of four different datasets: municipalities, MNIST, GLOVE and NYtimes. This study consists of a neighbours search applying Euclidean distance, Manhattan distance and Chebyshev distance. The main objective is to measure the accuracy of the GMASK algorithm with different distances and to observe if any distance is more adequate for some dataset, depending on its characteristics. For each dataset, we compare the results obtained by GMASK with two other algorithms: PYNN and FLANN. Both PYNN and GMASK are able to work with these three distances, however, FLANN does not work with Chebyshev distance. Since MASK is a specific case of GMASK only able to work with Euclidean distance, its inclusion in these experiments has not been considered.
Our approach to evaluating GMASK performance is conceived to closely emulate real-world conditions. The datasets used in this work are not artificially generated, and no preprocessing has been applied. The transformation and cleaning of the collected data is limited to dividing the original dataset into two distinct subgroups, the training and testing datasets. We also check that identical elements do not appear in both subgroups, thereby intensifying the complexity in the search for $k$ approximate neighbours.
@@ 4.1 Municipalities
The municipalities dataset is an interesting case due to the geographical distribution of points. In Figure 3, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands. The last two groups of points increase the difficulty of the analysis because they are far from the rest of the points, so we can consider them as outliers. Despite this dataset containing additional information, we only consider two variables: latitude and longitude.
Figure 4 shows the recall comparison between FLANN, PYNN and GMASK. We use a heatmap to represent the recall value obtained for each algorithm with each distance in the search of 5, 10 and 15 neighbours. The recall obtained in this case for each method is very similar, and all of them depict good recall values, over 84%. It is noticeable that FLANN recall decreases by about 10% with the number of neighbours. Regardless of the distance used and the number of neighbours intended to search, GMASK is the only method that provides a 100% recall in all tests.
Figure 4: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{muni}$.
@@ 4.2 MNIST
Given its high dimensionality and sparsity, MNIST has become a widespread benchmark in similarity search research. The following experiment evaluates the recall obtained in searching neighbours in this dataset, applying Manhattan, Euclidean and Chebyshev distance with FLANN, PYNN and GMASK methods. Figure 5 presents the results of this analysis.
From this comparison, we can deduce that the results are highly comparable in the recall when applying PYNN and GMASK methods with Euclidean and Manhattan distances, respectively. Therefore, the combination of both distances and these methods suits this case study. However, this is not the case for Chebyshev distance since the recall is lower than 70%. There is also a meaningful reduction in FLANN recall when the number of neighbours increases. Regardless of the distance, GMASK yields a higher average recall than other methods.
Figure 5: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{MNIST}$.
@@ 4.3 GLOVE
Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high dimensional datasets that contains the largest number of elements among the datasets included in these experiments.
Figure 6 depicts the recall results obtained for each method and distance. Once again, the luck of efficiency is evident in FLANN´s results. The best recall obtained with FLANN is 12% when searching for 5 neighbours applying Euclidean distance. PYNN method provides good accuracy when Manhattan distance is used. If we focus on PYNN method and compare the recall for the three distances, we can see that Manhattan distance provides the best results and Chebyshev distance provides the worst recalls. However, in the case of GMASK, the recall is 100% per each distance and with independence of the number of neighbours searched.
Figure 6: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{GLOVE}$.
@@ 4.4 NYtimes
Finally, the last experiment replicates the same analysis using a dataset characterised by high dimensionality, featuring a data-to-dimensionality ratio lower than that of GLOVE but higher than MNIST: the NYtimes dataset.
Figure 7 illustrates the recall accomplished by FLANN, PYNN and GMASK. There is a noticeable difference between GMASK and the other methods, since the recall values for FLANN and PYNN are below 50%. However, GMASK consistently outperforms the other methods regardless of the distance chosen. Once again, FLANN exhibits the poorest performance.
Therefore, while the most remarkable insight obtained from this experiment is the superior performance of the Chebyshev distance metric compared to others, it is crucial to highlight that, in this case, GMASK nearly doubles the recall provided by PYNN, achieving an 82% recall when searching for the 5 nearest neighbours.
Through this experiment, we empirically demonstrate what was previously stated theoretically in earlier sections of this manuscript: adopting the Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to sub-optimal results.
Figure 7: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{NYtimes}$.
@@ 5 Discussion
Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, which is strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems. Stemming from this need, GMASK makes the accommodation of arbitrary distances feasible by integrating the $k$-medoids algorithm as a clustering algorithm, allowing us to leverage the advantages derived from using several distance metrics on high-dimensional datasets.
It is crucial to emphasise that the algorithm’s generalisation, achieved through the integration of $k$-medoids, significantly broadens the range of applicable distances while preserving the structural integrity and properties of the original algorithm. Thus, GMASK maintains crucial characteristics, including the desirable scalability property facilitated by the bottom-top construction of the search tree, retaining a competitive edge over other state-of-the-art algorithms that require centralising all data on a single server.
As elaborated in Section “Experimental results”, our findings demonstrate the efficiency of this new approach, thereby supporting the hypothesis that alternative distances beyond the Euclidean metric may yield superior results in solving nearest-neighbour queries, contingent upon the specific characteristics of the dataset. Therefore, it becomes evident that the outcomes achieved using the proposed method surpass those obtained through two other widely adopted state-of-the-art algorithms, which are distinct from each other and significantly differ from GMASK.
@@
GMASK provides an avenue for future research to integrate distance functions beyond Minkowski, thus providing many opportunities to extend this research line. For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as cosine similarity for textual data or the Jaccard distance for categorical data, to address particular problems.
Furthermore, although GMASK has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance. To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PYNNDescent—, or considering the adaptation and port of GMASK to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.

### v2

@@ 3.4 Performance evaluation
A straightforward strategy to evaluate the performance of an algorithm is to search for the $k$ approximate nearest neighbours of every query point in a test dataset against a previously indexed train dataset.
To measure the accuracy of the algorithm’s performance in this task and thereby assess its ability to retrieve the relevant elements during a search, we would use the recall metric. This quality measure is defined as:
$Recall=\frac{\textit{Number of Relevant Items Retrieved}}{\textit{Total Numberof Relevant Items in the Search}}$
In this context, the Relevant Items are defined as the ground truth neighbours, corresponding to the true or correct set of $k$ points that the algorithm should ideally identify as the nearest neighbours of a given query point. They are obtained through the use of an exact search algorithm that will be chosen based to the distance metric employed, depending on its compatibility with the algorithm:
  • Brute-Force represents the most naive exact neighbour search implementation involves the computation of distances between all pairs of points in the dataset using a chosen metric (e.g. Euclidean distance, cosine similarity) and selecting the closest ones. Efficient brute-force neighbours searches can be very competitive for small data samples. However, as the number of samples grows, this approach quickly becomes infeasible.
  • The KD-Tree (K-Dimensional Tree) [64] serves as a widely used method for exact similarity search. It constructs a binary tree structure by recursively dividing the data based on the maximum spread along different dimensions and the search algorithm efficiently narrows down the search space by traversing the tree. In low-dimensional spaces, KD-trees can achieve sublinear search times. However, as the dimensionality increases, the performance of the KD-tree can degrade due to the Curse of Dimensionality phenomenon. Although it is the preferred algorithm for calculating ground truth neighbours due to its higher computational efficiency, it supports a smaller range of distance metrics compared to the Brute Force method.
@@ 3.5 Benchmark algorithms for similarity search
After defining our evaluation strategy, we shall compare the performance of GDASC in approximate nearest neighbour search against alternative MAM to validate our thesis:
  • FLANN (Fast Library for Approximate Nearest Neighbours) [15] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets. This library is extensively used in machine learning, computer vision and data mining. FLANN employs techniques such as random projection and hierarchical subdivision to construct an index structure that enables faster search operations. It offers several index types, including KD-tree, $k$-means tree and hierarchical clustering tree, which are suitable for diverse data types and query requirements. While FLANN has implementations in several programming languages, we use the Python bindings provided by the pyflann  [^2] https://github.com/primetang/pyflann library.
  • The PyNNDescent (PyNN) algorithm is an approximate similarity search method [37] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/. It is based on the NNDescent algorithm and is used for efficient nearest neighbour search in large-scale datasets. PyNN constructs a graph-based index by connecting each data point to its approximate nearest neighbours. It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching and pruning. The algorithm also provides parameters to control the trade-off between search accuracy and computational cost. PyNN is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining and information retrieval tasks.
@@ 3.6 Datasets
Several experiments have been conducted to evaluate the performance of GDASC using four publicly accessible datasets that vary in type, size, dimensionality (Dim), and sparsity (HS), with minimal or no preprocessing.Table 1 briefly summarises their main features.
Dataset Label n Dim HS Data Type Purpose
Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial Outliers
MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image High Dimensionality & Dispersion
GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text High Number of Elements
NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text Cosine Distance
Municipalities
@@
In Figure 5, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands. As the location of these islands deviates from the rest of the points as significantly, we can consider them as outliers.
Thus, despite being a two-dimensional dataset with a low dispersion, it is an exceptionally interesting case due to its intuitive representation and the presence of outliers, posing a significant challenge to any similarity search method applied.
MNIST
The MNIST_784 dataset  [^5] http://yann.lecun.com/exdb/mnist/ is a widely used machine learning benchmark dataset whose simplicity makes it an ideal starting point for algorithm development and evaluation. Firstly, it consists of 60,000 training images and 10,000 test images, a relatively small size that makes possible to carry out computationally efficient experimentation and prototyping. Secondly, it comprises grayscale images of handwritten digits ranging from 0 to 9, making it visually straightforward and easy to work with. Moreover, each image is 28x28 pixels in size, resulting in a total of 784 dimensions when flattened, allowing researchers to explore algorithms in the context of high-dimensional data. Many well-established techniques and algorithms have been evaluated on MNIST [65], making it a reference point for evaluating the performance of approximate similarity search methods [66, 49].
GLOVE
@@
NYtimes
The NYtimes dataset, part of the Bags of Words dataset from the UCI repository [^7] https://archive.ics.uci.edu/dataset/164/bag+of+words, comprises a collection of New York Times news articles represented as a bag of words. Each document in the dataset is associated with a set of word occurrences, where the dimensions represent unique words extracted from the articles. The dataset is organised as a document–word matrix, where each row corresponds to a document and each column corresponds to a word. The values in the matrix indicate the frequency of each word occurs in the respective document. Preprocessing steps include tokenization, removal of stopwords, and vocabulary truncation, with only words occurring more than ten times retained.
The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [67].
@@ 4 Experimental results
In order to prove the versatility of this framework, we have designed a series of experiments that may be repeated for every dataset described in Section 3.6. All of them have been randomly pre-split into train/test sets which, thereby intensifying the complexity, will not contain any common elements.
First, by using the train set, we approach MSA using Euclidean, Manhattan, Chebyshev, and cosine distances, respectively. Then, by using NSA we perform searches for the 5, 10, and 15 approximate nearest neighbours of every query point contained in the corresponding test set. To evaluate the performance of GDASC, we compute the average recall of the searches conducted using each distance metric and compare the results with those obtained by employing two other algorithms: PyNN and FLANN. Our framework enables the use of all the distances proposed for these experiments, similar to PyNN. However, FLANN does not support Chebyshev or cosine distances.
Once the average recall for each dataset-distance-method combination is obtained, we can determine if any specific distance metric is more suitable for a particular dataset based on its characteristics. This analysis supports one of the main hypotheses of this research: data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
@@
(a) $\mathcal{X}_{muni}$ dataset
(b) $\mathcal{X}_{MNIST}$ dataset
(c) $\mathcal{X}_{GLOVE}$ dataset
(d) $\mathcal{X}_{NYtimes}$ dataset
Figure 6: Point plot illustrating the Average Recall of three different algorithms across various distances (manhattan, euclidean, chebyshev, cosine) in approximate nearest neighbour search for each dataset
@@ 4.1 Municipalities
Figure 6a corresponds to a point plot that shows the average recall obtained in the $k$ approximate neighbours search with the FLANN, PyNN and GDASC methods, using Manhattan, Euclidean, Chebyshev and cosine distances. Regarding the GDASC algorithm, the structure generated by the MSA module has been built by setting the group length ($gl$) to 60 points, with 30 prototype points calculated for each group at the lowest level ($n_{1}$). Meanwhile, the NSA module uses a radius ($r$) of 30 to find the $k$ nearest neighbours within the structure created by the MSA module. For PYNNDescent and FLANN, the default parameter settings have been applied.
For this dataset, the average recall achieved using Minkowski distances is considerably high for all methods. In particular, PyNN and GDASC achieve almost 100% recall in these cases. However, while GDASC maintains near-perfect performance with the cosine distance, the average recall of PyNN drops dramatically to 54%, yielding the worst result for this dataset.
@@ 4.2 MNIST
MNIST is an interesting case of study due to its high dimensionality and sparsity. The results obtained in this experiment are presented in Figure 6b. Regarding the GDASC algorithm, the MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$, while NSA has a $r=80000$ settled. For PYNNDescent and FLANN, default parameter settings have been applied.
The analysis reveals that the average recall achieved by PyNN and GDASC are highly comparable, with GDASC consistently achieving slightly higher values. Both methods reach high recall values for all distances except for Chebyshev, which has an average recall below 70%. Additionally, the usefulness of FLANN is questioned not only by its lack of support for Chebyshev and cosine distances, but also by its low recall (less than 60%) for this dataset, which is more complex than the one analysed previously.
@@ 4.3 GLOVE
Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high-dimensional dataset that contains the largest number of elements among the datasets included in these experiments. Regarding GDASC, the MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$, while NSA is launched with $r=250$. As previously, default parameter settings have been applied for PYNNDescent and FLANN. Figure 6c depicts the average recall performed by each method and the distance evaluated.
Once again, the lack of efficiency is evident in FLANN´s results, which is not able to reach 15% with either the Euclidean or the Manhattan distance. If we focus on PyNN method and compare the recall for the four distances, we can see that the Manhattan distance provides a good accuracy and the Chebyshev distance gives the worst recall, while Euclidean and cosine distances provide a similar performance of over 60%. However, in the case of GDASC, it maintains a consistent average recall of 100%, regardless of the distance metric used.
@@ 4.4 NYtimes
Finally, the last experiment replicates the same analysis using a dataset characterised by high dimensionality, featuring a data-to-dimensionality ratio lower than that of GLOVE but higher than MNIST: the NYtimes dataset. In this scenario, GDASC MSA module is initialised with parameters $gl=1000$ and $n_{1}$$prototypes=500$ once again, while NSA is launched with $r=30$. For PYNNDescent and FLANN, default parameter settings have been applied.
Figure 6d illustrates the average recall achieved by FLANN, PyNN and GDASC, highlighting the most striking finding of this experiment: achieving good performance depends not only on the use of a suitable method, but also on the choice of an appropriate distance metric. Although GDASC consistently outperforms the other two algorithms in all scenarios, the trends observed for PyNN are similar to GDASC. Both algorithms exhibit the lowest recall with the Euclidean distance, followed by the Manhattan distance, and show considerable improvement with the Chebyshev distance. The best results are obtained with the cosine distance, where GDASC achieves an average recall close to 100%, significantly overtaking PyNN, which reaches an average recall slightly above 65%. Once again, FLANN exhibits the poorest performance.
Through this experiment, we empirically demonstrate what has been stated theoretically in earlier sections of this manuscript: adopting Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to suboptimal results.
@@ 5 Discussion
Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems[51].
Therefore, this paper introduces GDASC, a versatile data indexing framework that facilitates the incorporation of arbitrary distances by employing different clustering algorithms that assign a representative element to each cluster and allow the use of multiple $\ell_{p}$ norms or even other distance functions. This approach significantly broadens the spectrum of compatible problems while offering desirable scalability through the bottom-up construction of the search tree, providing a competitive edge over other state-of-the-art algorithms that require centralisation of all data on a single server.
As detailed in Section 4, experimental results using $k$-medoids clustering in GDASC with real datasets varying in type, size, dimensionality, and sparsity confirm the applicability of this approach for approximate similarity search, with outcomes surpassing those obtained with two other widely adopted MAM that differ significantly from GDASC. Our findings not only demonstrate the efficiency of this novel framework but also support the hypothesis that the use of alternative distances beyond the Euclidean metric may yield superior results for nearest-neighbour queries, depending on the specific characteristics of the dataset.
@@
GDASC provides an avenue for future research to integrate distance functions beyond Minkowski or cosine similarity, thus providing many opportunities to extend this research line. For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as the Jaccard distance for categorical data, to address particular problems.
Furthermore, although GDASC has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance. To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PyNN—, or considering the adaptation and port of GDASC to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.